OpenStudy (anonymous):

In a formula,z varies inversely as p. If z is 200 when p is 4, find z when p is 20.

OpenStudy (cwrw238):

z = k/p 200 = k/4 k = 200*4 = 800 so z = 800 / p now put p = 20 to find z

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

okay thanks :)

OpenStudy (cwrw238):

yw